

********************************************************************
UARS WEEKLY STATUS REPORT - 15 JUL 2005
********************************************************************
The decision regarding UARS by the 2005 Senior Review of Mission
Operations and Data Analysis Program for the Earth Sciences
Missions is summarized as follows (excerpt from letter from Mary
Cleave, Director, Earth-Sun System Division, to Charles
Jackman, UARS Project Scientist, dated July 7, 2005):
"You are to continue operating the UARS spacecraft and the HALOE
instrument until about January 1, 2006, at which time you are to
terminate the operations of the UARS spacecraft. All other
instrument operations are to cease as soon as possible upon
receipt of this letter."
Therefore, the operation of these UARS instruments will end on
August 1, 2005: PEM, HRDI, SUSIM, and SOLSTICE. Operation of
the HALOE instrument will continue, as spacecraft power permits,
through calendar year 2005.
***********
SOLSTICE
***********
Nothing received.
***********
SUSIM
***********
SUSIM was off this week. We anticipate a turn on scheduled for July 18.
***********
PEM
***********
HEPSA data files are version 2. Currently, HEPSA data between 1991274
and 2004273 has been reprocessed and archived at the GSFC DAAC. HEPSA data
between 2004274 and 2005079 has been processed and is waiting for archival
at the DAAC.
MEPS data files are version 1. Processed MEPS data files at the DAAC
archive are between 1991274 and 2004259. MEPS data between 2004260 and
2005069 has been processed and is waiting for DAAC archival. MEPS data for
2005070 to 2005079 is currently being processed.
PEM continues to receive NMC data files from NOAA. No problems have
been encountered since updates have occurred.
The PEM data from the GUVI conjunction periods may be found at
http://wwwpem.space.swri.edu/guvi-pem.html. The additional electron data for
the October and November storms of 2003 has been moved and is now located at
http://climatology.space.swri.edu/uars_fromRudy.html. For those users of
SDDAS, update your UARS Meta data and these data times will appear
with full functionality. Proton data from these storms are now located at
http://climatology.space.swri.edu/UARS_images_from_Rudy_(Apr._2004).html
Requests have been received for storm spectral data. Overview plots were
produced and these plots may be found on the web at
http://climatology.space.swri.edu/UARS_images_postedJune2004.html
for the storms of October 29 of 1991, November 8-9 of 1991, April 6-7 of 2000,
March 31 - April 1 of 2001, April 11-12 of 2001, November 6-7 of 2001, and
November 11-12 of 2001.
PEM has received a request for data from the super storm of November
2004. The Dst value showed sudden commencement early on November 7, 2004
reaching a Dst value of about -380 gamma early on November 8, 2004.
PEM was not active at that time and became active late on November 8, 2004
in day only mode. A second peak of about -280 gamma was observed on November
11, 2004. PEM became active in day-light only mode late on November 8. PEM
survey spectrograms for November 8-18 of 2004 may be found at:
http://climatology.space.swri.edu/UARS_images_fromRudy.html
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
MPG
***********
The MPG continues to serve as the point-of-contact for UARS
instrument operations in support of the Science Traceability
Mission of UARS.
Daily processing of instrument Daily Activity Plans continues
with no problems.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 02 July 2005 (Orbit 75572, GMT Day 183) through
08 July 2005 (Orbit 75677, GMT Day 189).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing)
and is performing nominally under two-battery operations. The
instrument operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
HALOE 186/2253 75632 POWER OFF
SUSIM 188/2205 75661 POWER OFF
SOLSTICE 188/2247 75662 FULLTIME TO DAY ONLY
The beta angle ranged between the angles of 34.2 degrees to 09.0
degrees for this report period. The beta angle is now decreasing
towards a minimum angle of 1 degree on 10 July 2005 (DOY 191).
Spacecraft battery 1, 2, and 3 performance monitoring continues.
Battery 1 remains OFF the charge relay. Battery 2 and Battery 3 are
currently maintaining greater than 23.6 V end-of-night (EON) load bus
voltage. Battery temperatures are stable with a temperature delta
between Battery 2 and Battery 3 of 4.15 to 3.81 degrees C. Additional
battery performance data for each battery on 01 July 2005 and 08 July
2005 are provided in Appendix D.
The clock error ranged from 10.3 to -10.0 msec during this report
period. Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT
for this report. A summary is listed below and the details of ALERT
and Out-Of-Limits occurrences are listed in Appendix B.
OUT-OF-LIMITS = 00
ALERTS = 00
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
FOT = 02
HALOE = 01
SUSIM = 01
SOLSTICE = 01
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
FOT = 02
Two (2) Anomaly Reports remained OPEN at the end of this report period
(see Appendix E).
GROUND SYSTEM OPERATION
Two (2) operational support problems occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations are no longer being supplied by Data
Capture Facility (DCF).
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
The power subsystem continued to be monitored closely last week. The
beta angle is now moving towards zero and instruments are being
powered off in preparation for the upcoming yaw on 10 July 2005.
During the low beta period the VT level will mostly likely have to be
raised to VT-6 to account for the reduced power availability and the
additional load during the yaw maneuver.
The UARS OBC clock rollover effort continued this week with more
discussions regarding contingency plans as well as an analysis of the
HGA slew angles during the uplink. The FOT also awaits the delivery
of a preliminary build for the Eclipse ground system, which is
expected sometime next week.
MEETINGS
The FOT attended a meeting with the PACOR software team on 06 July
2005 to discuss their readiness for the epoch year update. A software
fix is needed on their end to support the epoch update and progress of
this fix will be monitored via a weekly Wednesday meeting.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 10 July 2005 (DOY 191). This
maneuver will be a Forward to Reverse Flight maneuver (Northern
Hemisphere viewing).
Attachments: Appendix A, Real-time Command Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 08 July 2005
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
75632 186/2253 05-206 HALOE POWER OFF
75655 188/1229 05-207 OBC/CLK CLK ADJ FR 0.0 TO -20.6 MSEC/DAY
75661 188/2205 05-208 SUSIM SUSIM OFF
75662 188/2247 05-209 SOLSTICE SOLSTICE TO DAY ONLY
75677 189/2321 05-210 OBC/CLK CLK ADJ FR -20.6 TO 0.0 MSEC/DAY
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 08 July 2005
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period.
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period.
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 08 July 2005
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
75655 188/1240 CLK ADJ FR 0.0 TO -20.6 MSEC/DAY
75677 189/2323 CLK ADJ FR -20.6 TO 0.0 MSEC/DAY
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 08 July 2005
01 July 2005 - GMT Day 182 Beta = 35.9 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 11.2/ -11.2 -3.68/-4.94 +00.0/+00.0 05.8 0.00 0.0
2 -67.2/-246.4 +6.43/+5.14 +28.8/-18.0 25.8 1.03 21.4
3 123.2/-700.0 +3.23/+2.60 +21.6/-12.0 25.6 1.05 15.2
08 July 2005 - GMT Day 189 Beta = 09.0 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 22.4/ 0.0 -3.68/-4.94 +00.0/+00.0 05.8 0.00 0.0
2 0.0/-112.0 +6.44/+5.15 +31.2/-17.6 25.3 1.02 21.5
3 212.8/-700.0 +2.29/+1.34 +20.8/-09.6 25.1 1.04 13.4
V/T LEVEL CHANGES
ORB# AOS TIME LVL FR-TO
----- -------- ---------
-NONE THIS REPORT PERIOD
APPENDIX E
UARS ANOMALY REPORTS
Week ending 08 July 2005
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
05-032 FOT 75620 186/0356 30832 Generic Late Acquisition
05-033 FOT 75656 188/1303 40038 Late Acquisition
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
05-030 FOT See List June 05 30832 Generic Late Acquisition
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
05-032 FOT See List July 05 30832 Generic Late Acquisition
05-033 FOT 75656 188/1303 40038 Late Acquisition